OpenStudy (anonymous):

(____) ^2 = (secx -1) (secx +1) Help with this?

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

there is a standard identity \(\large \tan^2x=\sec^2x-1\) write 1 as 1^2 on right side,now apply \(a^2-b^2=(a+b)(a-b)\) what u get ?

OpenStudy (anonymous):

wait ...can you slow that down? 0.o

hartnn (hartnn):

remember this standard identity \(\large \tan^2x=\sec^2x-1\)

hartnn (hartnn):

ok ?

OpenStudy (anonymous):

Okay ..tan^2 x = sec^2 x -1 ..correct?

hartnn (hartnn):

now you see sec^2 x -1 ? write that as \(sec^2x-1^2\) now it is of the form \(a^2-b^2\) right ?

hartnn (hartnn):

?

OpenStudy (anonymous):

question tho.... it is that form but where is a2-b2 coming in at? 0.o

hartnn (hartnn):

i thought , u would never return from chat.....

OpenStudy (anonymous):

is that the secx + 1?

OpenStudy (anonymous):

my bad. well im done in chat now

OpenStudy (anonymous):

i seem to recall in algebra (x-2) (x+2) or somethign similar can be combined..right?

OpenStudy (anonymous):

thats what the right side part made me think of ..but i ended up getting confused..hence my question on here..

hartnn (hartnn):

\(sec^2x-1^2\) is in the form of \(a^2-b^2\) and \(a^2-b^2 = (a+b)(a-b)\) so, \(sec^2-1^2 = (sec x-1 )(secx+1)\) got this ?

OpenStudy (anonymous):

oh wait .... sec^2 -1^2 = (secx-1) (secx +1) right?

hartnn (hartnn):

thats what i told.

OpenStudy (anonymous):

yeah..its making sense now..but thne..thats the answer ? ...

hartnn (hartnn):

and since , u know \(\large \tan^2x=\sec^2x-1 \\ \large (\tan x)^2 = (secx-1)(sec x +1)\)

OpenStudy (anonymous):

oh ..that would be the identity ....so the answer is tan x?

hartnn (hartnn):

yes.

OpenStudy (anonymous):

i get that one too

hartnn (hartnn):

good :)

OpenStudy (anonymous):

might i make a suggestion that people stop just giving out answers...it doesn't really help :/

OpenStudy (anonymous):

they need to do it like you ...

hartnn (hartnn):

you can express your suggestions/feedback in 'Openstudy Feedback' Group.....